Use the related acute angle to state an equivalent expression:
a) tan 110 degrees
b)sin 350 degrees
Answer #1
a) tan 110 degrees = tan 110-180=
tan-70
b)sin 350 degrees = sin 350-360
= sin-10 degrees
answered by: Henry P
Answer #2
The more standard practise is to state the equivalent angle in terms of a positive acute angle.
So ...

tan 110° = -tan 70°
sin 350° = -sin 10°
answered by: HELP! :)